# Robust Multi-Modal Density Estimation

### v1

@@ 4 Experiments
as well as using $k_{\min}=5$, $k_{\max}=20$, and $\alpha_{k}=400$ for calculating $k_{\bm{c}}$ (Equation (2)).
@@ 4.2 Evaluation and Metrics
  • To test for over-fitting, we first sample two different datasets $\bm{X}_{1}$ and $\bm{X}_{2}$ ($N$ samples each) from $p$ ($\bm{X}_{1},\bm{X}_{2}\sim p$). We then use the estimator $f$ to create two queryable distributions $\widehat{p}_{1}=f(\bm{X}_{1})$ and $\widehat{p}_{2}=f(\bm{X}_{2})$. If those distributions $\widehat{p}_{1}$ and $\widehat{p}_{2}$ are similar, it would mean the tested estimator does not over-fit; we measure this similarity using the Jensen–Shannon Divergence Lin (1991):
@@
  • Not every density estimator $f$ has the ability to generate the samples $\widehat{\bm{X}}_{1}$. Consequently, we need to test for over-smoothing (and over-fitting as well) without relying on $\widehat{\bm{X}}_{1}$. Therefore, we use the average log-likelihood
@@
    which would be maximized only for $p=\widehat{p}_{1}$ as long as $\bm{X}_{2}$ is truly representative of $p$ (see Gibbs’ inequality Kvalseth (1997)). However, using this metric might be meaningless if $\widehat{p}_{1}$ is not normalized, as the presence of the unknown scaling factor makes the $\widehat{L}$ values of two estimators incomparable.
@@ 4.3 Ablations
  • Decorrelation vs No decorrelation. We investigated the effect of ablating rotation by setting $\bm{R}_{C}=\bm{I}$.
@@
    where $\mathcal{V}_{M}$ is the volume of the $M$-dimensional unit hypersphere. We used the rule-of-thumb $k=\lfloor\sqrt{N}\rfloor$. However, this estimator does not enable sample generation.
@@ 5.1 Baseline Comparison
Clustering
Decorr. No decorr.
Silhouette
$13\scriptscriptstyle{\pm 0.20}- 0$
$13\scriptscriptstyle{\pm 0.20}- 0$
$13\scriptscriptstyle{\pm 0.19}- 0$
$14\scriptscriptstyle{\pm 0.20}- 0$
DBCV
$09\scriptscriptstyle{\pm 0.23}- 0$
$09\scriptscriptstyle{\pm 0.23}- 0$
$07\scriptscriptstyle{\pm 0.23}- 0$
@@
Clustering
Decorr. No decorr.
Silhouette
$-2.53\scriptscriptstyle{\pm 0.02}$
$-2.70\scriptscriptstyle{\pm 0.01}$
$-2.79\scriptscriptstyle{\pm 0.01}$
DBCV
$-2.56\scriptscriptstyle{\pm 0.02}$
$-2.69\scriptscriptstyle{\pm 0.01}$
$-2.83\scriptscriptstyle{\pm 0.02}$
We found that ROME avoids major pitfalls of the two baseline methods on the four tested distributions (Table 1). Out of the two baseline methods, the Manifold Parzen Windows approach has a stronger tendency to over-fit in the case of the two-dimensional distributions compared to our approach, as quantified by lower $D_{\text{JS}}$ values achieved by ROME. MPW does, however, achieve a better log-likelihood for the Two Moons distribution compared to ROME. This could be due to the locally adaptive non-Gaussian distributions being less susceptible to over-smoothing than our approach of using a single isotropic kernel for each cluster if such clusters are highly non-normal. Lastly, in the case of the pedestrian trajectories distribution, our approach once more achieves better performance than MPW, with MPW performing worse both in terms of $D_{\text{JS}}$ and the log-likelihood estimates.
Meanwhile, the Vine Copulas approach exhibits the tendency to over-smooth the estimated densities (large positive $\widehat{W}$ values), and even struggles with capturing the different modes (see Figure 4). This is likely because VC uses KDE with Silverman’s rule of thumb, which is known to lead to over-smoothing in the case of multi-modal distributions Heidenreich et al. (2013). Furthermore, on the pedestrian trajectories distribution, we observed both high $D_{\text{JS}}$ and $\widehat{W}$ values, indicating that VC is unable to estimate the underlying density; this is also indicated by the poor log-likelihood estimates.
@@
Norm. No norm.
$f_{\text{GMM}}$
Clustering
Decorr. No decorr.
Silhouette
${1.40}\scriptscriptstyle{\pm 0.52}$
$1.41\scriptscriptstyle{\pm 0.52}$
$3.65\scriptscriptstyle{\pm 0.99}$
$4.37\scriptscriptstyle{\pm 1.14}$
DBCV
$1.59\scriptscriptstyle{\pm 0.56}$
$1.59\scriptscriptstyle{\pm 0.56}$
$4.24\scriptscriptstyle{\pm 1.13}$
$4.77\scriptscriptstyle{\pm 1.23}$
No clusters
$1.82\scriptscriptstyle{\pm 0.60}$
$1.84\scriptscriptstyle{\pm 0.60}$
$3.17\scriptscriptstyle{\pm 0.89}$
$5.29\scriptscriptstyle{\pm 1.34}$
Refer to captionROMERefer to captionKDE
@@
Normalization No normalization
$f_{\text{GMM}}$
@@
Clustering
$f_{\text{KDE}}$
$f_{\text{kNN}}$
$f_{\text{KDE}}$
$f_{\text{kNN}}$
$f_{\text{KDE}}$
$f_{\text{kNN}}$
Silhouette
$0.084\scriptscriptstyle{\pm 0.016}$
$1.045\scriptscriptstyle{\pm 0.064}$
$0.777\scriptscriptstyle{\pm 0.116}$
$1.808\scriptscriptstyle{\pm 0.112}$
$0.015\scriptscriptstyle{\pm 0.011}$
$1.887\scriptscriptstyle{\pm 0.118}$
$0.032\scriptscriptstyle{\pm 0.007}$
DBCV
$0.090\scriptscriptstyle{\pm 0.015}$
$1.119\scriptscriptstyle{\pm 0.073}$
$0.897\scriptscriptstyle{\pm 0.154}$
$1.937\scriptscriptstyle{\pm 0.109}$
$0.017\scriptscriptstyle{\pm 0.012}$
$1.934\scriptscriptstyle{\pm 0.116}$
$0.043\scriptscriptstyle{\pm 0.010}$
No clusters
$0.009\scriptscriptstyle{\pm 0.004}$
$0.453\scriptscriptstyle{\pm 0.051}$
$0.015\scriptscriptstyle{\pm 0.012}$
$1.044\scriptscriptstyle{\pm 0.104}$
$0.005\scriptscriptstyle{\pm 0.003}$
$1.478\scriptscriptstyle{\pm 0.132}$
$0.017\scriptscriptstyle{\pm 0.011}$
@@ 5.2 Ablation Studies
Testing variants of ROME on the Aniso distribution (Table 3) demonstrated not only the need for decorrelation, through the use of rotation, but also normalization in the case of distributions with highly correlated features. There, using either of the two clustering methods in combination with normalization and decorrelation (our full proposed method) is better than the two alternatives of omitting only decorrelation or both decorrelation and normalization. In the case of clustering with the silhouette score, the full method is significantly more likely to reproduce the underlying distribution $p$ by a factor of $1.19$ ($p<10^{-50}$) as opposed to omitting only decorrelation, and by $1.30$ compared to omitting both decorrelation and normalization. Similar trends can be seen when clustering based on DBCV, with the full method being more likely to reproduce $p$ by a factor of $1.14$ and $1.31$ respectively. Results on the Aniso distribution further show that KDE on its own is not able to achieve the same results as ROME, but rather it has a tendency to over-smooth (Figure 5).
Additionally, the ablation with and without normalization on the Two Moons distribution (Table 4) showed that normalization is necessary to avoid over-smoothing in the case of non-normal distributions.
Lastly, investigating the effect of different downstream density estimators, we found that using ROME with $f_{\text{kNN}}$ instead of $f_{\text{KDE}}$ leads to over-fitting (highest $D_{\text{JS}}$ values in Table 5). Meanwhile, ROME with $f_{\text{GMM}}$ exhibits a tendency to over-smooth the estimated density in cases where the underlying distribution is not Gaussian (high $\widehat{W}$ in Table 4). The over-smoothing caused by $f_{\text{GMM}}$ is further visualised in Figure 6.
In conclusion, our ablation studies confirmed that using $f_{\text{KDE}}$ in combination with data clustering, normalization and decorrelation provides the most reliable density estimation for different types of distributions without relying on prior knowledge.
Refer to captionROMERefer to captionROME with $f_{\text{GMM}}$
@@ 6 Conclusion
In our comparison against two established and sophisticated density estimators, we observed that ROME achieved consistently good performance across all test cases, while Manifold Parzen Windows and Vine Copulas were susceptible to over-fitting and over-smoothing. Furthermore, as part of several ablation studies, we found that ROME could overcome the shortcomings of other common density estimators, such as the over-fitting exhibited by kNN or the over-smoothing by GMM. In those studies, we additionally demonstrated that our approach of using clustering, decorrelation, and normalization is indispensable for overcoming the deficiencies of KDE. Future work can further improve on our results by investigating the integration of more sophisticated density estimation methods, such as MPW or VC, instead of the kernel density estimator in our proposed approach.
Overall, by providing a simple way to accurately estimate learned distributions based on model-generated samples, ROME can help evaluate AI models in non-deterministic scenarios, paving the way for more reliable AI applications across various domains.

### v2

@@ 4 Experiments
as well as using $k_{\min}=5$, $k_{\max}=20$, and $\alpha_{k}=400$ for calculating $k_{\bm{c}}$ (Equation (5) and Appendix C).
@@ 4.1 Distributions
Further tests on uni-modal problems can be found in Appendix E.
@@ 4.2 Evaluation and Metrics
  • To test for over-fitting, we first sample two different datasets $\bm{X}_{1}$ and $\bm{X}_{2}$ ($N$ samples each) from $p$ ($\bm{X}_{1},\bm{X}_{2}\sim p$). We then use the estimator $f$ to create two queryable distributions $\widehat{p}_{1}=f(\bm{X}_{1})$ and $\widehat{p}_{2}=f(\bm{X}_{2})$. If those distributions $\widehat{p}_{1}$ and $\widehat{p}_{2}$ are similar, it would mean the tested estimator does not over-fit; we measure this similarity using the Jensen-Shannon divergence Lin (1991):
@@
    This metric, however, is not able to account for systematic biases that could be present in the estimator $f$. Comparisons of $\widehat{p}_{1}$ with $p$ – if the dataset allows – can be found in Appendix D.
@@
  • Not every density estimator $f$ has the ability to generate the samples $\widehat{\bm{X}}_{1}$. Consequently, we need to test for goodness-of-fit without relying on $\widehat{\bm{X}}_{1}$. Therefore, we use the average log-likelihood
@@
    which would be maximized only for $p=\widehat{p}_{1}$ as long as $\bm{X}_{2}$ is truly representative of $p$ (see Gibbs’ inequality Kvalseth (1997)). However, using this metric might be meaningless if $\widehat{p}_{1}$ is not normalized, as the presence of the unknown scaling factor makes the $\widehat{L}$ values of two estimators incomparable, and cannot discriminate between over-fitting and over-smoothing.
@@ 4.3 Ablations
  • Decorrelation vs No decorrelation. We investigated the effect of removing rotation by setting $\bm{R}_{C}=\bm{I}$.
@@
    where $\mathcal{V}_{M}$ is the volume of the $M$-dimensional unit hypersphere. We used the rule-of-thumb $k=\lfloor\sqrt{N}\rfloor$. However, this estimator cannot generate samples.
@@ 5.1 Baseline Comparison
Clustering Decorr. No decorr.
Silhouette $13\scriptscriptstyle{\pm 0.20}- 0$ $13\scriptscriptstyle{\pm 0.20}- 0$ $13\scriptscriptstyle{\pm 0.19}- 0$ $14\scriptscriptstyle{\pm 0.20}- 0$
@@
Clustering Decorr. No decorr.
Silhouette $-2.53\scriptscriptstyle{\pm 0.02}$ $-2.70\scriptscriptstyle{\pm 0.01}$ $-2.79\scriptscriptstyle{\pm 0.01}$
DBCV $-2.56\scriptscriptstyle{\pm 0.02}$ $-2.69\scriptscriptstyle{\pm 0.01}$ $-2.83\scriptscriptstyle{\pm 0.02}$
We found that ROME avoids the major pitfalls displayed by the two baseline methods on the four tested distributions (Table 1). Out of the two baseline methods, the manifold Parzen windows (MPW) approach has a stronger tendency to over-fit in the case of the two-dimensional distributions compared to ROME, as quantified by lower $D_{\text{JS}}$ values achieved by ROME. MPW does, however, achieve a better log-likelihood for the Two Moons distribution compared to ROME. This could be due to the locally adaptive non-Gaussian distributions being less susceptible to over-smoothing than our approach of using a single isotropic kernel for each cluster if such clusters are highly non-normal. Lastly, in the case of the pedestrian trajectories distribution, ROME once more achieves better performance than MPW, with MPW performing worse both in terms of $D_{\text{JS}}$ and $\widehat{L}$.
Meanwhile, the vine copulas (VC) approach tends to over-smooth the estimated densities (large positive $\widehat{W}$ values), and even struggles with capturing the different modes (see Figure 4). This is likely because VC uses KDE with Silverman’s rule of thumb, which is known to lead to over-smoothing in the case of multi-modal distributions Heidenreich et al. (2013). Furthermore, on the pedestrian trajectories distribution, we observed both high $D_{\text{JS}}$ and $\widehat{W}$ values, indicating that VC is unable to estimate the underlying density; this is also indicated by the poor log-likelihood estimates.
@@
Norm. No norm. $f_{\text{GMM}}$
Clustering Decorr. No decorr.
Silhouette ${1.40}\scriptscriptstyle{\pm 0.52}$ $1.41\scriptscriptstyle{\pm 0.52}$ $3.65\scriptscriptstyle{\pm 0.99}$ $4.37\scriptscriptstyle{\pm 1.14}$
DBCV $1.59\scriptscriptstyle{\pm 0.56}$ $1.59\scriptscriptstyle{\pm 0.56}$ $4.24\scriptscriptstyle{\pm 1.13}$ $4.77\scriptscriptstyle{\pm 1.23}$
No clusters $1.82\scriptscriptstyle{\pm 0.60}$ $1.84\scriptscriptstyle{\pm 0.60}$ $3.17\scriptscriptstyle{\pm 0.89}$ $5.29\scriptscriptstyle{\pm 1.34}$
Refer to captionROME $f_{\text{GMM}}$efer to captionKDE $f_{\text{GMM}}$
@@
Normalization No normalization $f_{\text{GMM}}$
@@
Clustering $f_{\text{KDE}}$ $f_{\text{kNN}}$ $f_{\text{KDE}}$ $f_{\text{kNN}}$ $f_{\text{KDE}}$ $f_{\text{kNN}}$
Silhouette $0.084\scriptscriptstyle{\pm 0.016}$ $1.045\scriptscriptstyle{\pm 0.064}$ $0.777\scriptscriptstyle{\pm 0.116}$ $1.808\scriptscriptstyle{\pm 0.112}$ $0.015\scriptscriptstyle{\pm 0.011}$ $1.887\scriptscriptstyle{\pm 0.118}$ $0.032\scriptscriptstyle{\pm 0.007}$
DBCV $0.090\scriptscriptstyle{\pm 0.015}$ $1.119\scriptscriptstyle{\pm 0.073}$ $0.897\scriptscriptstyle{\pm 0.154}$ $1.937\scriptscriptstyle{\pm 0.109}$ $0.017\scriptscriptstyle{\pm 0.012}$ $1.934\scriptscriptstyle{\pm 0.116}$ $0.043\scriptscriptstyle{\pm 0.010}$
No clusters $0.009\scriptscriptstyle{\pm 0.004}$ $0.453\scriptscriptstyle{\pm 0.051}$ $0.015\scriptscriptstyle{\pm 0.012}$ $1.044\scriptscriptstyle{\pm 0.104}$ $0.005\scriptscriptstyle{\pm 0.003}$ $1.478\scriptscriptstyle{\pm 0.132}$ $0.017\scriptscriptstyle{\pm 0.011}$
@@ 5.2 Ablation Studies
Testing variants of ROME on the Aniso distribution (Table 3) demonstrated not only the need for decorrelation, through the use of rotation, but also normalization in the case of distributions with highly correlated features. There, using either of the two clustering methods in combination with normalization and decorrelation (our full proposed method) is better than the two alternatives of omitting only decorrelation or both decorrelation and normalization. In the case of clustering with the silhouette score, the full method is significantly more likely to reproduce the underlying distribution $p$ by a factor of $1.19$ (with a statistical significance of $10^{-50}$, see Appendix B) as opposed to omitting only decorrelation, and by $1.30$ compared to omitting both decorrelation and normalization. Similar trends can be seen when clustering based on DBCV, with the full method being more likely to reproduce $p$ by a factor of $1.14$ and $1.31$ respectively. Results on the Aniso distribution further show that KDE on its own is not able to achieve the same results as ROME, but rather it has a tendency to over-smooth (Figure 5). Additionally, the ablation with and without normalization on the Two Moons distribution (Table 4) showed that normalization is necessary to avoid over-smoothing in the case of non-normal distributions.
Lastly, investigating the effect of different downstream density estimators, we found that using ROME with $f_{\text{kNN}}$ instead of $f_{\text{KDE}}$ leads to over-fitting (highest $D_{\text{JS}}$ values in Table 5). Meanwhile, ROME with $f_{\text{GMM}}$ tends to over-smooth the estimated density in cases where the underlying distribution is not Gaussian (high $\widehat{W}$ in Table 4). The over-smoothing caused by $f_{\text{GMM}}$ is further visualised in Figure 6.
In conclusion, our ablation studies confirmed that using $f_{\text{KDE}}$ in combination with data clustering, normalization and decorrelation provides the most reliable density estimation for different types of distributions.
Refer to captionROME $f_{\text{GMM}}$efer to captionROME with $f_{\text{GMM}}$
@@ 6 Conclusion
In our comparison against two established and sophisticated density estimators, we observed that ROME achieved consistently good performance across all test cases, while manifold Parzen windows (MPW) and vine copulas (VC) were susceptible to over-fitting and over-smoothing. For example, while MPW is superior at capturing non-normal distributions, it produces kernels with too small of a bandwidth (hence the over-fitting), which is likely caused by the selected number of nearest neighbors used for the localized kernel estimation being too small. Meanwhile, compared to VC, ROME is numerically more stable and does not hallucinate new modes in the estimated densities (Figure 4). Furthermore, as part of several ablation studies, we found that ROME overcomes the shortcomings of other common density estimators, such as the over-fitting exhibited by kNN or the over-smoothing by GMM. In those studies, we additionally demonstrated that our approach of using clustering, decorrelation, and normalization is indispensable for overcoming the deficiencies of KDE.
Future work can further improve on our results by investigating the integration of more sophisticated density estimation methods, such as MPW, instead of the kernel density estimator in our proposed approach to enable better performance on non-normal clusters.
Overall, by providing a simple way to accurately estimate distributions based on samples, ROME can help in better handling and evaluating probabilistic data as well as enabling more precise probabilistic inference.